AMadge invested $2,500 at 8% interest for one year. She has an additional $8,000 to deposit. At what rate must she invest the $8,000 to end the year with a total balance of exactly $11,540?
Answer by mananth(16946)

AMadge invested $2,500 at 8% interest for one year. She has an additional $8,000 to deposit. At what rate must she invest the $8,000 to end the year with a total balance of exactly $11,540?
$2,500 at 8%
8000@x%
Balance = interest + investment
investment = 8000+2500 = 10500
Balance = 11540
so she has to earn an interest of 11540-10500=1040
Interest equation
2500*8%+8000*x%= 1040
200+80x=1040
80x=1040-200
80x=840
x=840/80
x=10.5%